a) Determine the mean molar mass of the atmosphere of Venus, which consists of 95% CO2 and 5% N2 by volume.


b) What is the corresponding gas constant?


c) The mean surface temperature T on Venus is a scorching 740K as compared to only 288K for Earth; the surface pressure is 90 times that on Earth. By what factor is the density of the near-surface Venusian atmosphere greater or less than that of Earth?


1
Expert's answer
2020-09-08T05:56:57-0400

a) Mean molar mass is

M = 0.95*44 (CO2) + 0.05*28 (N2) = 43.2 g/mole

b) gas constant R = 8.314 J/K.mole

specific gas constant is

Rspec = R/M = 8.313/43.2 = 0.192 J/K.g

c) from gas law

p*V = m*R*T/M and density d = m/V we have

p*M/(R*T) = d (here R = 8.314 J/K.mole)

on earth molar mass of air is 29.0 g/mole

Hence, on Venus the atmosphere is density by the factor:

d(Venus)/d(Earth) = p(Venus)/p(Earth)*M(Venus)/M(Earth)*T(Earth)/T(Venus) = 90*43.2/29.0*288/740 = 52.17

answer: by the factor of 52.17